Product Pricing using the Cost-Plus Approach Methods; Differential Analysis for Accepting Additional Business

Crystal Displays Inc. recently began production of a new product, flat panel displays, which required the investment of $1,500,000 in assets. The costs of producing and selling 5,000 units of flat panel displays are estimated as follows:

Variable costs per unit: Fixed costs:
Direct materials $120 Factory overhead $250,000
Direct labor 30 Selling and administrative expenses 150,000
Factory overhead 50
Selling and administrative expenses 35
Total variable cost per unit $235

Crystal Displays Inc. is currently considering establishing a selling price for flat panel displays. The president of Crystal Displays has decided to use the cost-plus approach to product pricing and has indicated that the displays must earn a 15% return on invested assets.

Required:

Note: Round all markup percentages to two decimal places, if required. Round all costs per unit and selling prices per unit to the nearest whole dollar.

1. Determine the amount of desired profit from the production and sale of flat panel displays.
$

2. Assuming that the product cost method is used, determine the following:

a. Product cost amount per unit $
b. Markup percentage %
c. Selling price per unit $

3. (Appendix) Assuming that the total cost method is used, determine the following:

a. Total cost amount per unit $
b. Markup percentage %
c. Selling price per unit $

4. (Appendix) Assuming that the variable cost method is used, determine the following:

a. Variable cost amount per unit $
b. Markup percentage %
c. Selling price per unit $

5. The cost-plus approach price computed above should be viewed as a general guideline for establishing long-run normal prices; however, other considerations, such as the price of competing products and general economic conditions of the marketplace , could lead management to establish a different short-run price.

6. Assume that as of August 1, 3,000 units of flat panel displays have been produced and sold during the current year. Analysis of the domestic market indicates that 2,000 additional units are expected to be sold during the remainder of the year at the normal product price determined under the product cost method. On August 3, Crystal Displays Inc. received an offer from Maple Leaf Visual Inc. for 800 units of flat panel displays at $225 each. Maple Leaf Visual Inc. will market the units in Canada under its own brand name, and no variable selling and administrative expenses associated with the sale will be incurred by Crystal Displays Inc. The additional business is not expected to affect the domestic sales of flat panel displays, and the additional units could be produced using existing factory, selling, and administrative capacity.

a. Prepare a differential analysis of the proposed sale to Maple Leaf Visual Inc. If an amount is zero, enter "0".

Differential Analysis
Reject (Alt. 1) or Accept (Alt. 2) Order
August 3
Reject
Order
(Alternative 1)
Accept
Order
(Alternative 2)
Differential
Effects
(Alternative 2)
Revenues $ $ $
Costs
Variable manufacturing costs
Profit (loss) $ $ $

Feedback
Answer #1

I am posting answer for the first 4 parts of the question

1. The amount of desired profit from the production and sale of flat panel displays

The president of crystal displays has decided to use the cost plus approach to product pricing and indicated that the diaplays must earn a 15% return on invested assets

Given, Investments in assets = 1,500,000

Hence the desired profit = Investment * return

= 1.50,000*0.15 = 225,000

2. Assuming the Product cost method,

a) Product cost amount per unit

Product cost does not include sales and administrative expenses both variable and fixed.

Cost amount per unit = (Direct material + Direct labour + Variable Factory overhead) per unit + Fixed factory overhead per unit

= (120+30+ 50)+ (250.000/5000)

  = (200)+(50)

  = 250 per unit

b) Markup Percentage

Markup cost = Desired profit / Number of units

= 225000/5000

  \large = $45 per unit

Markup Percentage = Markup / Cost amount per unit.

  - 45/250

  = 0.180r18%

Selling price per unit = Variable cost per unit + Fixed cost per. unit + Markup

= 235 + [(250,000+150,000)/5000] + 45

= 235 + 80 + 45

= $ 360 per unit

3. Assuming the total cost method,

a) Cost amount per unit = Variable cost per unit + fixed cost per unit

    = 235 + [(250,000 + 150,000)/5000]

= 235 + 80

= $ 315 per unit

b) Markup Percentage

Desired profit = $ 225,000

Units = 5000 units

Mark up = 225000/5000

= $ 45 per unit

Mark up percentage = Markup / Cost amount per unit

= 45/315

= 0.143 or 14% (round off)

c) Selling price per unit

= Cost + Markup

= 315 + 45

= $360 per unit

4. Assuming Variable cost method

  a) Variable cost amount per unit

Variable cost is given in the question, = $235 per unit

b) Markup percentage

Markup percentage = Markup / Variable cost

= 45 / 235

= 0.191 or 19%(round off)

c) Selling price per unit

Selling price = Total cost + Markup

= 315 + 45

= $ 360 per unit